OpenStudy (anonymous):

Candace said, "I'm thinking of a fraction. if i divide it by 1/2, i get 3/11." what fraction was candace thinking of?

OpenStudy (anonymous):

do the reverse take 3/11 and multiply by 1/2 and get 3/22... that is what candace was thinking of

OpenStudy (anonymous):

3/22 ?

OpenStudy (anonymous):

is the answer

OpenStudy (anonymous):

x/1/2 =3/11 2x =3/11 x = 3/22

OpenStudy (anonymous):

Yupp.